Given: 1/3(one third)x + y = 15.


Which of the following are ordered pairs of this equation?

A.{(-3, 14),(0, 15),(3, 16)}
B.{(-3, 16),(0, 15),(3, 14)}
C.{(16, -3),(15, 0),(14, 3)}

Answer by ReadingBoosters(3246)

Starting with
1/3x + y = 15
Put it into the linear equation format y = mx+b
y = -1/3x +15
...
Take -3, 0 and 3 as x and solve for y
y = -1/3(-3) + 15 = 16: -3,16
y = -1/3(0) + 15 = 15: 0, 15
y = -1/3(3) + 15 = 14: 3, 14
...
Answer B
............
Delighted to help.
-Reading Boosters
Wanting for other what we want for ourselves.
www.MyHomeworkAnswers.com